OpenStudy (anonymous):

Given: F(x) = 2x - 1; G(x) = 3x + 2; H(x) = x 2 Find F(x) + G(x) + H(x).

OpenStudy (anonymous):

Given: F(x) = 2x - 1; G(x) = 3x + 2; H(x) = x 2 Find F[G(x)] - F(x).

OpenStudy (ajprincess):

jst add the three functions. F(x) + G(x) + H(x).= 2x - 1+3x + 2+ x 2 Combine the like terms. Does that help?

OpenStudy (anonymous):

@ajprincess i figure out that one

OpenStudy (anonymous):

5x+1+x^2

OpenStudy (ajprincess):

to find f(g(x))(x) u have to plug in g(x) in pace of x in f(x). f(g(x))(x)=2g(x) - 1=2( 3x + 2)-1 f(g(x))(x)-f(x)=2( 3x + 2)-1-( 2x - 1)

OpenStudy (anonymous):

4x

OpenStudy (ajprincess):

i guess the answer 5x+1+x^2 is for the first ques right? If so then it is right:)

OpenStudy (anonymous):

it wasn't a answer choice

OpenStudy (ajprincess):

well what r ur answer choices?

OpenStudy (anonymous):

4x + 4 4x + 2 4x

OpenStudy (ajprincess):

u got 5x+1+x^2 for which question?

OpenStudy (anonymous):

nvm , i got it

OpenStudy (anonymous):

Given: A = {(0, 0), (2, 1), (3, 1.5), (4, 2), . . . } Set A is a function.

OpenStudy (ajprincess):

f(g(x))(x)-f(x)=2( 3x + 2)-1-( 2x - 1)=6x+4-1-2x+1=4x+4

OpenStudy (anonymous):

Given: A = {(0, 0), (2, 1), (3, 1.5), (4, 2), . . . } Set A is a function.

OpenStudy (anonymous):

true or false

OpenStudy (ajprincess):

sorry no idea.:(